# Physics beyond the Standard Model with the DSA-2000

### v1

@@ 6 Discussion and conclusions
Science case Data Time Resolution Sensitivity ($g_{a\gamma\gamma}/\epsilon$) Mass ($\mu$eV)
Axion DM high-res beam $10$ h $10$ kHz $4\times 10^{-14}$ GeV-1 $2.9-8.3$
Axion clouds imaging/ high-res beam $10$ h/ $4$ h $50$ MHz $4\times 10^{-15}$ GeV-1 $2.9-8.3$
@@
Table 6: Table summarizing requirements and resulting sensitivity for three of the considered science cases with electromagnetic signatures. The imaging data product is the standard continuum all-sky survey image cube. High-resolution beamforming data refers to placing a single beam on the position of a known source, preserving data down to 100 $\mu$s sampling time and 10 kHz.
One significant finding concerns axion cloud formation around NSs, where we have developed an analytical treatment of the expected radio signal as a function of period and magnetic field, previously examined only through simulations. Our analytical framework conservatively captures key features of the full simulation while providing physical insight into the underlying processes. Using targeted observations not yet included in the planned observing strategy, DSA-2000 could potentially reach sensitivity to the theoretically significant QCD axion target band in the mass range 2.9--8.3 $\mu$eV. This represents a promising avenue for detecting axions, with potential implications for axion direct detection.
@@
  • Axion miniclusters have observable radio signals. However, the prediction for axion miniclusters constituting all of the DM lies above the DSA-2000 frequency band [38, 40, 39, 255, 256].
  • Stimulated axion decay [257] triggered by sources in the sky such as remnant supernova, lead to radio emission. The signal is most competitive for frequencies below the DSA-2000 band.
  • The decay or scattering of weakly interacting massive particles into leptons induces a radio signal via synchrotron radiation, if the leptons are accelerated by astrophysical magnetic fields [258, 259, 260, 261, 262]. This promising signal is sensitive to the astrophysical environment.
  • Collisions between DM in the form of ultraheavy composite bound states, so-called dark blobs, could induce transient radio signals [263] which DSA-2000 is optimized for, a promising target to pursue in future work.
  • The large number of FRBs projected to be detected by DSA-2000 lead to a small chance of measuring a repeating lensed FRB. Such a lucky measurement would provide a very sensitive probe of DM substructure [264]. Unfortunately, due to the scanning strategy of DSA-2000, and the theoretical gaps in being able to predict in which interval an FRB may repeat, make this an extremely challenging measurement.
@@ Acknowledgments
We are grateful to Masha Baryakhtar, Joshua Foster, David E. Kaplan, Andrea Mitridate, Vikram Ravi, Benjamin Safdi, Kai Schmitz, Myles Sherman, Erwin Tanin and Tanner Trickle for stimulating discussion. We are especially grateful to Joseph Lazio, Cristina Mondino, Stephen Taylor and Samuel Witte for providing comments on the manuscript. K.B. thanks the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632 and the Walter Burke Institute for Theoretical Physics. V.L. is supported by NSF Physics Frontier Center Award #2020275 and the Heising-Simons Foundation. A.P. acknowledges support from the Princeton Center for Theoretical Science. K.Z. is supported by a Simons Investigator Award, the U.S. Department of Energy, Office of High Energy Physics under Award Number DE-SC0011632, and by the Walter Burke Institute for Theoretical Physics.

### v2

@@ 6 Discussion and conclusions
[…]
Axion DM high-res beam $10$ h $10$ kHz $2\times 10^{-13}$ GeV-1 2.98.32.9-8.3
Axion clouds imaging/ high-res beam $10$ h/ $4$ h $50$ MHz $4\times 10^{-15}$ GeV-1 2.98.32.9-8.3
@@
[…]
@@
  • Axion miniclusters have observable radio signals. However, the prediction for axion miniclusters constituting all of the DM lies above the DSA-2000 frequency band [Gorghetto:2018myk, Buschmann:2019icd, Gorghetto:2020qws, Edwards:2020afl, Gorghetto:2024vnp].
  • Stimulated axion decay [Buen-Abad:2021qvj] triggered by sources in the sky such as remnant supernova, lead to radio emission. The signal is most competitive for frequencies below the DSA-2000 band.
  • The decay or scattering of weakly interacting massive particles into leptons induces a radio signal via synchrotron radiation, if the leptons are accelerated by astrophysical magnetic fields [Colafrancesco:2005ji, 2013ApJ...768..106S, Leite:2016lsv, Regis:2021glv, Bhattacharjee:2020phk]. This promising signal is sensitive to the astrophysical environment.
  • Collisions between DM in the form of ultraheavy composite bound states, so-called dark blobs, could induce transient radio signals [Kaplan:2024dsn] which DSA-2000 is optimized for, a promising target to pursue in future work.
  • The large number of FRBs projected to be detected by DSA-2000 lead to a small chance of measuring a repeating lensed FRB. Such a lucky measurement would provide a very sensitive probe of DM substructure [Xiao:2024qay]. Unfortunately, due to the scanning strategy of DSA-2000, and the theoretical gaps in being able to predict in which interval an FRB may repeat, make this an extremely challenging measurement.
@@ Acknowledgments
We are grateful to Masha Baryakhtar, Abhiram Cherukupalli, Joshua Foster, David E. Kaplan, Andrea Mitridate, Vikram Ravi, Benjamin Safdi, Kai Schmitz, Myles Sherman, Erwin Tanin and Tanner Trickle for stimulating discussion. We are especially grateful to Joseph Lazio, Cristina Mondino, Stephen Taylor and Samuel Witte for providing comments on the manuscript. K.B. and Y.D. thank the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632 and the Walter Burke Institute for Theoretical Physics. V.L. is supported by NSF Physics Frontier Center Award #2020275 and the Heising-Simons Foundation. A.P. acknowledges support from the Princeton Center for Theoretical Science. K.Z. is supported by a Simons Investigator Award, the U.S. Department of Energy, Office of High Energy Physics under Award Number DE-SC0011632, and by the Walter Burke Institute for Theoretical Physics.